OpenStudy (anonymous):

a) Consider the reaction CaCO3(s) <- -> CaO(s) + CO2(g). Should this experiment be performed in an open or closed container if I want to use up the most amount of the reactant? Explain. b) Does an exothermic reaction go more to completion if I raise the temperature, lower the temperature or perform the experiment at room temperature? Explain. c) What color would I expect a solution of chromate ions to be after I add an excess of acid? Explain. d) What color would I expect a solution of dichromate ions to be after I add an excess of base? Explain. e) Which one of the following equations would make more product upon addition of acid (H+)? Explain. 1) H2SO4(aq) <--> HSO4-(aq) + H+(aq) 2) HCl (aq) <--> H+(aq) + Cl-(aq) 3) NH4+(aq) + OH-(aq) <--> NH3(aq) + H2O(l) 4) 2Fe2O3 (s) <--> 2FeO(s) + O2(g) 5) H+(aq) + NH3(l) <--> NH4+(aq) Any help would be greatly appreciated.

OpenStudy (anonymous):

a) cao doesn't react directly with co2..... first you have to make calcium hydroxide by reactiong it with water.. cao + h2o = ca(oh)2 then ca(oh)2+co2 = caco3 + h2o... first dissolve cao in water... never mind to close it.... then purge the co2 gas in to the liquid (ca(oh)2) .. here also no need to close it.... cause our main component is calcium hydroxide which is in liquid state..... so you will get it totally reacted......

OpenStudy (anonymous):

thanks @harishk for the help!

OpenStudy (anonymous):

c) What color would I expect a solution of chromate ions to be after I add an excess of acid? Explain. d) What color would I expect a solution of dichromate ions to be after I add an excess of base? Explain. would that shift from yellow to orange and orange to yellow?